OpenStudy (katherinesmith):

Which of the following is the equation of g(x) if the parent function is f(x) = x^2 and a = 3, h = -5, and k = 6?

OpenStudy (katherinesmith):

f(x) = 3(x - 5)^2 + 6 f(x) = -3(x + 5)^2 + 6 f(x) = 3(x + 5)^2 + 6 f(x) = 3(x - 5)^2 - 6

OpenStudy (anonymous):

general equation for parabola: p(x) = a(x-h)^2 + k

OpenStudy (katherinesmith):

so C?

OpenStudy (anonymous):

yes

OpenStudy (katherinesmith):

thank you!